MAIK.BST

1. \refitem{techreport } BSI, BS 1629, British Standards Institution (1976).

2. \refitem{techreport } BSI, BS 5606, British Standards Institution (1978).

3. \refitem{techreport } BSI, BS 6371, British Standards Institution (1983).

4. \refitem{book} J. Butcher, Copy-editing (Cambridge University Press, 1981), 2nd ed.

5. \refitem{book} ‘Chicago’, The Chicago Manual of Style (University of Chicago Press, 1982), 13th ed.

6. \refitem{article } W. W. Howells, American Journal of Physical Anthropology 9, 159 (1951).

7. \refitem{article } W. W. Howells, Current Anthropology 7, 531 (1966).

8. \refitem{article } W. W. Howells, Annals of the New York Academy of Sciences 134, 624 (1966).

9. \refitem{article } G. B. Johnson, Science 184, 28 (1974).

10. \refitem{article } G. B. Johnson and W. W. Howells, Journal journal journal (1974).

11. \refitem{article } G. B. Johnson, W. W. Howells, and A. N. Other, Journal journal journal (1976).

12. \refitem{manual } Anon., Title title title title title title title title title title, Organization organization organization (1967).

13. \refitem{article } E. J. Aslin, Aslib Proceedings 1, 49 (1949).

14. \refitem{article } W. M. Fletcher and F. G. Hopkins, J. Physiol. 35, 247 (1907).

15. \refitem{incollection } C. W. Hanson, in Handbook of special librarianship and information work, edited by W. Ashworth (1967), pp. 414–452, 3rd ed.

16. \refitem{inbook } R. C. Wright, Report Literature (1963), pp. 46–59.

17. \refitem{book} Association of British Chemical Manufacturers, British chemicals and their manufacturers (1959).

18. \refitem{inbook } F. Feigl, Spot Tests in Organic Analysis (Publisher publisher, 1958), chap. 6, 5th ed.

19. \refitem{book} P. G. Hoel, Elementary Statistics, Wiley series in probability and mathematical statistics (Wiley, New York, Chichester, 1971), 3rd ed., iSBN 0 471 40300.

20. \refitem{inbook } P. G. Hoel, Elementary Statistics (Wiley, New York, Chichester, 1971), pp. 19–33, Wiley series in probability and mathematical statistics, 3rd ed., iSBN 0 471 40300.

21. \refitem{book} C. J. Singer, E. J. Holmyard, and A. R. Hall, eds., A history of technology (Oxford University Press, London, 1954–58), 5 vol.

22. \refitem{incollection } in A history of technology, edited by C. J. Singer, E. J. Holmyard, and A. R. Hall (Oxford University Press, London, 1954–58), vol. 5.

23. \refitem{incollection } in A history of technology, edited by C. J. Singer, E. J. Holmyard, and A. R. Hall (Oxford University Press, London, 1954–58), vol. 5.

24. \refitem{manual } BSI, Natural Fibre Twines, British Standards Institution, London, 3rd ed. (1973), bS 2570.

25. \refitem{techreport }

BSI, BS 2570, British Standards Institution, London (1973), 3rd. edn.

26. \refitem{inbook } BSI, BS 2570: Natural Fibre Twines (British Standards Institution, London, 1973), Table 5, 3rd ed.

27. \refitem{techreport } B. Ellis and A. K. Walton, Tech. Rep. RAE-TR-71009, Royal Aircraft Establishment (1971).

28. \refitem{article } G. B. Godfrey, Struct. Eng. 37, 126 (1959).

29. \refitem{incollection } J. Ramsbottom, in A System of Bacteriology in relation to Medicine (HMSO, for Medical Research Council, London, 1931), vol. 8, pp. 11–70.

30. \refitem{article } J. Hanlon, New Scientist pp. 429–432 (1972).

31. \refitem{techreport } Winget Ltd., GB Patent Specification 1060631 (1967).

32. \refitem{article } I. Bry and L. Afflerbach, Community Mental Health 4, 75 (1968).

33. \refitem{incollection } S. R. Ranganthan, in Bibliographic Organization, edited by J. H. Shera and M. E. Egan (1951), pp. 94–105.

34. \refitem{book} L. R. McColvin, Libraries in Britain (Longmans Green, for the British Council, London, ????).

35. \refitem{unpublished } Exchequer (1634–1639), edinburgh, Scottish Record Office, E.4/5.

36. \refitem{unpublished } E. Traquair (1638), lennoxlove (E. Lothian), Muniments of Duke of Hamilton and Brandon, C.1, no. 963.

37. \refitem{unpublished } J. Pym (1624), northampton, Northamptonshire Record Office, Finch-Hatton 50.

38. \refitem{phdthesis }

W. B. Croft, Ph.D. thesis, Cambridge University (1978).

39. \refitem{book} F. D’Arcy, Title title title title title title title title title title (Publisher publisher publisher, 1920).

40. \refitem{book} F. D’Arcy, Title title title title title title title title title title (Publisher publisher publisher, 1920).

41. \refitem{article } F. Jones, F. Norman, F. Hazel, and F. Robinson, Journal journal journal (1962).

42. \refitem{article } F. Jones, F. Smith, and F. Robinson, Journal journal journal (1962).

43. \refitem{article } F. Jones, Journal journal journal (1965).

44. \refitem{article } F. Jones, Journal journal journal (1969).

45. \refitem{article } F. Jones and F. Abrams, Journal journal journal (1968).

46. \refitem{article } F. Jones and F. Smith, Journal journal journal (1965).

47. \refitem{article } F. Jones, F. Abrams, and F. Smith, Journal journal journal (1966).

48. \refitem{article } A. J. Bloggs, Journal journal journal (1960).

49. \refitem{article } A. J. Bloggs, Journal journal journal (1961).

50. \refitem{article } A. J. Bloggs and X. Y. Jones, Journal journal journal (1959).

51. \refitem{article } A. J. Bloggs, R. S. Smith, and X. Y. Jones, Journal journal journal (1955).

52. \refitem{article } H. W. Brown, A. S. Forbes, and S. D. Smith, Journal journal journal (1900).

53. \refitem{incollection } P. Eckstein and S. Zuckermann, in Marshall’s Physiology of Reproduction, edited by A. S. Parkes (Longman, London, 1960), vol. 1, pp. 43–154.

54. \refitem{article } H. Heller and K. Lederis, Nature 182, 1231 (1958).

55. \refitem{book} R. H. Wood, Plastic and Elastic Design of Slabs and Plates (Thames & Hudson, London, 1961).

56. \refitem{book} F. Pratt, Title title title title title title title title title title (Publisher publisher publisher, 1975).

57. \refitem{article } F. Light, Journal journal journal (1972).

58. \refitem{article } F. Light and F. Wong, Journal journal journal (1975).

59. \refitem{article } F. Kingston, Journal journal journal (1976).

60. \refitem{article } F. Kelley, Journal journal journal (1896).

61. \refitem{article } F. Kelley, Journal journal journal (1896).

62. \refitem{article } F. Kelley, Journal journal journal (1907).

63. \refitem{book} F. Strong, Title title title title title title title title title title (Publisher publisher publisher, 1901).

64. \refitem{book} F. Strong, Title title title title title title title title title title (Publisher publisher publisher, 1902).

65. \refitem{book} F. Shotwell, Title title title title title title title title title title (Publisher publisher pub-

lisher, 1901).

66. \refitem{techreport } ‘Brunswick’, Tech. Rep. 1984, Rodent Activities Termination Section (RATS), Pest Control Division, Brunswick Public Welfare Department, Hamelin (1985).

67. \refitem{book} S. C. Kendeigh, Parental care and its evolution in birds, vol. 22 of Illinois Biological Monographs (Univ. of Illinois Press, Champaign, 1952).

68. \refitem{book} J. Chapman, The Icehouse Bottom Site—40MR23, no. 23 in University of Tennessee Department of Anthropology Publication (Univ. of Tennessee Press, Knoxville, 1975).

69. \refitem{book} P. Hershkovitz, Evolution of Neotropical cricetine rodents (Muridae) with special reference to the phyllotine group, vol. 46 of Fieldiana: Zoology (Field Museum of Natural History, Chicago, 1962).

70. \refitem{book} S. Wright, Evolution and the genetics of populations, vol. 4 (Univ. of Chicago Press, Chicago, 1978).

71. \refitem{incollection } S. Wright, in Evolution and the genetics of populations (Univ. of Chicago Press, Chicago, 1978), vol. 4.

72. \refitem{incollection } D. Ogilvy, in The Creative Organization, edited by G. A. Steiner (University of Chicago Press, Chicago, 1965), pp. 199–213.

73. \refitem{incollection } W. H. McNeill, in The Rise of the West (University of Chicago Press, Chicago, 1963), part 1.

74. \refitem{incollection } V. Thomson, in American Music since 1910 (Holt, Rinehart and Winston, New York, 1971), chap. 8.

75. \refitem{incollection } R. Gordon, in Playwrights for Tomorrow: A Collection of Plays, edited by A. H. Ballet (University of Minnesota Press, Minneapolis, 1975), vol. 13, one of four plays included in vol. 13.

76. \refitem{incollection } J. Milton, in The Complete Poetical Works of John Milton, edited by W. V. Moody (Houghton Mifflin, Boston, 1924), Student’s Cambridge ed.

77. \refitem{inproceedings } N. Chomsky, in A festschrift for Morris Halle, edited by S. R. Anderson and P. Kiparsky (Holt, Rinehart & Winston, New York, 1973).

78. \refitem{inproceedings } K. E. Chave, in Approaches to paleoecology, edited by J. Imbrie and N. Newel (Wiley, New York, 1964), pp. 377–87.

79. \refitem{book} N. Smart, The religious experience of mankind (Schribner, New York, 1976), 2nd ed.

80. \refitem{article } R. Jackson, Australian Geographer 14, 175 (1979).

81. \refitem{article } O. Prufer, Scientific American pp. 90–102 (1964).

82. \refitem{phdthesis } A. E. Mann, Ph.d. diss., University of California, Berkeley (1968).

83. \refitem{phdthesis } J. Maguire, Ph.d. diss., Department of Geology, University of the Witwatersrand, Johannesburg (1976).

84. \refitem{unpublished } W. J. Downes (1974), london School of Economics. Mimeo.

85. \refitem{misc } H. [pseud.] Hunt, Interview, Tape recording, Pennsylvania Historical and Museum Commission, Harrisburg (1976), interview by Ronald Schatz, 16 May 1976.